TCCON > Sites > East Trout Lake > Instrument History > Site Visit 2016-09-19 - 2016-09-30

Site Visit 2016-09-19 - 2016-09-30

To purchase

Hardware

  • RONA Hardware list 
  • duct tape
  • aluminum tape
  • surface pressure funnel
  • cable wrap
  • vacuum grease
  • lens paper & clamp
  • squeeze bottle
  • alcohol
  • kim wipes

Electronics

  • assortment of wire

To do

  1. Set up Cryomech system; generate at least one tank for the Norhof system (OC)
  2. Install Zeno Weather Station; hook up to Herc (JM)
  3. Install Herc and ensure that it can be reached from outside the lab; Senen to port forward 4222 on his router to 192.168.1.87 (our router's IP). Our router is already set up to port forward its port 22 to the herc. (DW/Senen)
    1. Move webcam onto router; change port to 8080 (port forward 8080 to the webcam IP, which should be changed to 10.10.0.5); gateway 10.10.0.52
    2. Put IFS125HR webpage through router; port 80 (port forward 80 to the IFS125HR, which should already be set to 10.10.0.1)
    3. Not sure what to do about the suntracker, whether we should port forward another http port to 10.10.0.2 (which should already be the IP of the tracker). Maybe port 4280 for the tracker.
  4. Cut foam to fit IFS125HR (after Gregor is finished assembly) (OC)
  5. Hook up micropirani gauge to herc; and pump to pump switch box; hook up fiber between switch box and herc
  6. Set up pump; attach to IFS125HR (GS?)
  7. Connect Norhof system to InSb detector
  8. Fabricate heater/temperature sensor cable to reach suntracker (OC)
  9. Compare Zeno pressure/temperature/RH to ECCC's weather station
  10. Take picture of inside of pump switch box; attach to wiki
  11. Strip paint off IFS125HR for heaters
  12. Attach temperature sensors to IFS125HR
  13. Check Norhof temp sensor cables to see if 2m one exsits (longer than 1.25 m which was shipped)
  14. O2 sensor installed and connected to the herc.
  15. Mount Measurement Computing modules somewhere near herc.
  16. Measure altitude of solar tracker mirrors once installed: Distance between IFS and first tracker mirror is ~3m.

 

September 19, 2016

  • Travel day. Gregor, Ivan, Orfeo and Joseph arrived at the site.
  • Crates were delivered and unloaded, no apparent damage.
  • Raining so deferred opening crates until tomorrow.

September 20, 2016

  • Opened IFS125HR crates; removed sample compartment and excess resolution compartment.
  • Scanner ribbon cable channel is too long; will need to be cut down.
  • InGaAs detector found inside the IFS125HR! Will be returned to Bruker for a refund.
  • Webcam moved to router. Webcam and herc can be accessed from inside router, but not yet externally. Hopefully Senen can help us solve this problem by tomorrow morning.
  • Weather station installed on the roof, but not yet connected through the wall of the lab.
  • We will move the Cryomech LN2 generator into the lab first thing tomorrow morning, then the IFS125HR can be put together.

 weatherstation_rot.jpeg

September 21, 2016

  • Lots of progress today!

IFS125HR

  • In the morning, Gregor and Ivan put together the Bruker, installed the laser, and by lunch were ready to turn everything on.
  • The voltage in the lab was a bit high (~125VAC), so the ETON UPS came in handy. The instrument, laser and suntracker will be plugged in there.
  • After lunch, they systematically turned the system on: all voltages were okay, laser stabilized as expected.
  • All filter wheels turned, all mirrors slid, all pinhole wheels rotated.
    • Note: we have a filter wheel in front of the InGaAs and InSb detector, so we will use that for the InSb longpass filter instead of our complicated post/holder system. We're trying to decide whether to disconnect the filter wheel power so it never turns, or just leave the filter in the default position, so it always moves there automatically after a reboot.
  • Aligned the laser.
    • Gregor aligned the laser in a clever way: in front short adjust mode, he placed a card in front of LASB(?-the one *not* on the same side of the beamsplitter as the input window (which, incidentally, is a red filter to prevent light leakage from other wavelengths entering the IFS)) and coaligned the two laser spots using the prism adjustment screws: the one from the fixed and moving cube corners. He then sent the scanner to back short adjust mode and ensured that the two laser spots were again coaligned.
    • He then deliberately separated the two laser points so that he could adjust the beam expander focus and ensured that the two beam spots were about the same size and ~3mm in diameter. Once they were about the same size, he coaligned them again.
    • The laser amplitudes did not perceptibly change as a function of OPD. (I did not measure the ME.)
  • Hooked up the pump; pumping down overnight.
  • We realized that the InGaAs detector I purchased and modified was not set up for dual DC detection. The original InGaAs was, in fact, in the IFS all along, so we will have to modify the original InGaAs to have the test points pulled out for future alignment.
  • The InSb does not have dual DC capabilities, but Gregor knew that ahead of time and is prepared to modify the preamp accordingly.
  • The highest phase resolution we can use on this instrument is 0.2cm-1.

LN2 Generation System

  • Everyone helped move the LN2 system into the back room.
  • Orfeo and Stephen set up the air compressor and purged it for 1.5 hrs in the afternoon. After purging it was noticed the flow rate was very low. Orfeo decided to increase the pressure in the LN2 dewar and purge the system again. 
  • After purging Orfeo and Stephen went to turn on the Cryomech LNP-10 He compressor and found that the Electrical protection deviceowed was in fault mode. Opening up the cover showed that the ABB CM-EFS voltage monitoring relay had a over/under voltage warning. This was most probably due to the high voltage in thelab 242 VAC as opposed to the 208-230 VAC the LNP-10 required. This high range had been approved earlier with the Cryomech technician Brian Stoddard. Orfeo has emailed Brian to re-confirm and see if we can alter the range which the Electrical protection device is tripping on.We did not started generating LN2 due to this fault.
  • After ~3 hrs of operation of the air compressor and N2 generation the O2 sensor began to trip it's alarm with levels dropping to 19.7%  O2. The air compressor was turned off and a leak was found on the Nitrogen flex line entering the dewar. We left the lab for dinner with the assumption that O2 levels would rise with the air compressor turned off. After 1.5 hours the O2 level has risen to 20.0 %

Norhof Micro-Doser

  • Orfeo, Joseph and Debra set up the microdoser system. We will test it out when we have some LN2!
  • The Norhof system only has ~1.5 m dosing line, much less than other systems. We plan to use the current line and then retrieve a longer one from Toronto this weekend. 

Network

  • Senen helped Debra set up the router so it would port forward the webcam. Turns out that the gateway on the webcam was set incorrectly, and after that was fixed, the problem was resolved. The webcam can be accessed by pointing your browser to etl.ftpaccess.cc:8080

Miscellaneous

  • Joseph and Debra built a tool chest. It was not very Ikea. But it's done, now.

Tomorrow

  • Tomorrow, we'll tackle the tracker and the LN2 line leak.

Back to top

September 22, 2016

  •  Clear skies. Time to couple the sun tracker to the Bruker.

Sun Tracker

  • Gregor and Ivan spent most of the morning fighting with Java to run the suntracker program. A full set of instructions on setting up JAVA on Window7 is attached.
  • Gregor and Ivan spent most of the day aligning the sun tracker with the Bruker.
  • Gregor said it was achieving 0.25o accuracy. Which according to Gregor is really good considering the modifications to the tracker.
  • Orfeo and Joseph put the enclosure back on the sun tracker and reconnected the enclosure motor.
  • The cable routing plate for the sun tracker cables was also installed. The heater cables for the suntracker cover (SIL heat pads and HERC heat pads) are not yet installed.
  • The suntracker was blocked by the vertical post on the north-west side of the railing which holds the Environment Canada pyranometer. This was blocked between 17:13-17:18. W . I belive this puts the sun around 8 Elevation, 260 azimuth.
  • Measured the distance between the first suntracker mirror and the 45-degree mirror as about 3 m: 273 cm between the plate holding Gregor's bob and the 45-degree mirror, and a few cm between the plate and the first tracker mirror.

Zeno

  • Orfeo mounted the Zeno box to the wall next to the pump.
  • The Zeno box is now connected to the Herc but no sensors are wired in yet.
  • Orfeo drilled a hole in the container and installed the pipe so that the sensors can be wired to the Zeno. 
  • We still need to re-oganise the order of the variables which are logged by the zeno

O2 Sensor

  • O2 sensor was calibrated.
  • Orfeo and Joseph connected the O2 sensor (which is in the LN2 generator room) to the Herc.

Herc

  • Debra and Joseph moved the Herc to the desk next to the pump.
  • It was reconnected to the network and is now accessible.
  • The micropirani was connected to the pump and the Herc.
  • The Zeno and O2 sensor are connected to the Herc but not tested

LN2

  • Orfeo eventually got in contact with Brian Stoddard from Cryomech and was able to confirm that the Electrical protection device could have it's overvoltage threshold increaed to allow the higher 242 VAC to run on the LNP-10. Brian said to adjust the top trim pot on the ABB CM-EFS up slightly to allow the higher voltage. Orfeo turned the top >U trim pot clock wise to increase the value from 255 VAC to 270 VAC. 
    • note that the spec sheet for the ABB CM-EFS says the top trim pot adjust the over voltage hysterisis. Perhaps the initial power up of the He compressor draws more that 242 VAC and increasing the hysterisis alows this to settle. 
  • This adjustment cleared the fault mode from the electrical protection device.
  • The initial stats on the LNP-10 are He Low pressure 70-80 PSI, He High pressure 280-285 PSI, dewar pressure 7.8 psi, filter bank 85 psi. Average time between air compressor re-charging 8-9 min. After running for 7 hors the LN2 level was 3/32 full (~3.3L). We hope there'll be enough LN2 to cool the detectors tomorrow. 
  • The O2 sensor was not tripped by the LN2 generation. The level was between 20.6-20.9%. 

Bruker 125HR

  • Gregor noticed that the filter wheel in the source compartment was a little bent from it's transport in the accessories box. The bent portion was at the larger diameter apertures. Gregor has removed the aperture and swapped it to the exit aperture position as this aperture is less critical. Gregor is not worried by the warping as it is not in the position of the smaller diameter that we use. 

Tomorrow

  • We will connect the sensors to the Zeno.
  • Gregor will install the InSb detector (with out modification) and run and N2O cell measurements while trying to aligning the instrument. 

September 23, 2016

  • A rainy day

IFS125HR

  • Gregor and Ivan ran cell tests the entire day in order to characterize the ILS. Gregor used his N2O and HCl cells as well as the ETL HCL cell (only 10 scans).
  • Joseph analyzed the spectrum using LINEFIT. The analysis shows an over modulation, which Gregor says is consistent with what they found at Bruker.

Zeno

  • Orfeo, Joseph, and Stephen hooked up the sensors to the Zeno logger.
  • The Zeno was attached to the Herc using the fiber.

LN2

  • Orfeo showed Stephen and Joseph the daily checks for the LN2 generator.

  • Joseph did the daily checks for the LN2 system at the end of the day. The Dewar is 1/3 full.

Tomorrow

  • Analyze HBr cell test.
  • Gregor will try to improve alignment.
  • Analyze cell tests.
  • Mark where heat pads and temp sensors will go on IFS.

September 24, 2016

  • Another rainy day

IFS125HR

  • Gregor used LINEFIT9 to analyze the HBr cell spectra that were measured over night. The mod eff @ max OPD (65 cm) was found to be 0.985.
  • Gregor aquired another HCl cell spectrum using the ETL HCl cell but this time co-adding 50 scans (HCl spectrum attached).
  • Joseph used LINEFIT12 to analyze the spectum. @ max OPD (45cm) the mod eff was found to be 1.061.
  • Debra used LINEFIT14.5 on the same spectrum and found the mod eff to be ~1.01. The difference between the analysis done by Joseph and Debra (besides the version of LINEFIT) was that Debra used the effective pressure (@ 296K) measured by Frank Hase while Joseph caculated the pressure from the temperature (294.4K, this was the temperature of the interferometer compartment) and column amount in the cell (measured by Frank Hase). When Joseph used the same pressure as Debra in LINEFIT12 he got the same results. See attached file ETL_HCL_ILS_Test Sept24,2016_JM.pdf.
  • Gregor modified the InGaAs and InSb detectors to work together. InSb was modefied using the notes from Jean-Francois.   
  • Joseph stripped the paint off the spots where the temperture sensors will go on the Bruker.

LN2

  • Joseph did the daily checks for the LN2 generator. All indicators show nominal operation.
  • The tank is 1/3 full.

Water leaking through sun tracker pipe.

  • Joseph found drops of water on the photodiode tube. Ivan and Joseph checked the suntracker to find the leak. Water does not apper to be leaking into the sun tracker. water could be leaking from the where the pipe is attached to the roof. 

September 25,2016

  • Windy and Sunny.

Update on leak

  • Joseph, Gregor, and Ivan inspected the suntracker and found that no water had entered the enclosure. The only place it could have come from is black plate is attached to the pipe that enters the building.

IFS125HR

  • Gregor is monitoring how much the tracker drifts throughout the day while tracking the sun.
  • Internal battery on tracker motor controller died. This caused the date to reset to 1980 when the controller was plugged in. Ivan went to Prince Albert, got a replacement battery, and replaced it.
  • GPS that came with the sun tracker is not functioning well. Only worked for first half of day.
  • Joseph stripped the paint from the Bruker for the heat pads.

Zeno

  • Norton was still unable to communicate with the Zeno today.
  • Joseph was able to communicate with the Zeno via hyperterminal and the RS-232 cable that came with it. No communication with the Zeno was possible using the fiber. Multiple set up were tried but none worked.
  • Joseph changed the records in the Zeno to match the record order (found at http://sol.caltech.edu/Devices/Weath...utrecords.html). He flipped the record postions of: the barometer and Solar sensor, the leaf wet and Precip sensor, and the Bit and battery. The output string should now match the caltech string order. Attached are screen shots of the settings in the Data ouput menu of the Zeno before they were changed by Joseph (Zeno Default Records settings in Data output menu.docx).
  • Norton was able to communicate with the Zeno once the fiber was removed and the RS-232 cable was plugged into the Herc.

LN2

  • Joseph did the daily checks. Indicators are normal, except high and low pressure on the cryomech are the same values (~200 psi) Probably because  dewar is full.

September 26, 2016

  • Another sunny day

IFS125HR

  • The pump has been pummping down the system to 0.5 mbar the last two days but when it was first turned on it could reach 0.2 mbar. Gregor connected his gauge directly to the pump to measure it and found the same value. Can't explain the drop in performance over a short period of time. 

Zeno

  • Jean-Francois remembered that the hardware handshake on the Zeno RS-232 cable had to be disabled inorder to use the fiber. To do this one has to go connect to the Zeno and type: u <Enter>                    [ To get the User menu ]
                  c                   [ Communications menu ]
                  m                   [ Modem menu ]
                  c1/0                [ was "1" as shipped from manufacturer ]
                  c2/NONE             [ was "COM3" as shipped from manufacturer ]
                  e                   [ save to EEPROM ]
  • Norton disabled the hardware handshake. Joseph connected to the Zeno using hyperterminal and the Zeno RS-232 cable and found that he could still talk to the Zeno.
  • Debra suggested that the Zeno RS-232 null-modem cable should not be connected to the fiber and the fiber should be connected directly to the Zeno and herc. Orfeo remembered that this is not possible since the nuts on the adapters get in the way (confirmed by Joseph) and suggested using serial cables along with the fiber. Joseph removed the Zeno RS-232 null-modem cable and tried to connect to the Zeno using the fiber with serial cables using hyperterminal but this did not work either. He tried  switching between dte and dce but neither worked. Also, tried powering both of the 9271A's as well as not powering the 9271A connected to the Zeno. None of these worked. 

 

 Back to top

September 27, 2016

Zeno

  • The only way we were able to get the Zeno talking to Hyperterminal is the following:

    • Original Zeno RS-232 null modem cable (we believe) to the Telebyte driver set to DTE. The fiber is connected to the DTE driver in the following manner: B2 on the Receive (R) port, A1 on the Transmit (T) port. On the second driver, set to DCE, B1 is connected to R, A2 is connected to T.

    • Unfortunately, to connect to the herc, we need another cable, because the drivers are too large to fit behind the herc and plug in, so I made a 2-3-5 straight through cable.
    • Also, this doesn't let you get into the menus. I think I will need to make a full straight-through cable. Found a serial cable and that did the trick.
    • Herc is connected to the Zeno with the optical fiber.

InSb

  • To set the gain on the InSb detector, it must be addressed individually first, set to REF gain, and then it can be addressed in combination with InGaAs. 
  • 50-50 beamsplitter installed between InSb and InGaAs. Looks okay; running lamp scans overnight.

Pump

  • The pump is now unable to achieve a pressure below 1.5 hPa when connected to the instrument, or much below 0.75 hPa when a blank is directly attached to the pump itself. This was not the case last night. Not sure of the cause. Perhaps a leak in the tubing to the micropirani (which wasn't touched) or the pump is on its last legs. It has not be used before, but did sit in a box for ~5 years.

Temperature Sensors and Heaters

  • The temperature sensors are hooked up and glued to the IFS (except the electronics (taped for now), Zeno (not glued or taped yet), laser (taped)).
    • We checked that the temperatures changed correctly on the IFSloop when we warmed up the individual sensors.
  • The heaters are ready to go; paint has been removed from the IFS, and cable is ready, just no access to the IFS to stick them down.
  • Orfeo is working on the tracker heater and temp sensor cable.

Norhof

  • Set up and ready to be tested tomorrow.

Enclosure

  • Norton fixed a couple of bugs, but the enclosure is opening and closing on command
    • Enclosure close
    • Enclosure open

O2 Sensor

  • Norton happy with the O2 sensor driver; need to test the alarm with nitrogen.

NTP

  • Need to figure out how to port forward the NTP servers to the herc.

Tomorrow

  • NTP
  • tracker heater/temp sensor cable
  • connect IFS and tracker to herc
  • install heaters on IFS
  • test O2 sensor alarm
  • look for leaks in pump
  • ...

Back to top

September 28, 2016

  • Port forwarded UDP port 123 to 10.10.0.8 (the herc). NTP now working.
  • Tested O2 sensor alarm by putting LN2 near the sensor
  • Pump cleaned, helped initial pumpdown.
  • Tried to let IFSloop take control of all hardware
  • Suntracker: Gregor unhappy with alignment. In TPS the solar image drifts off the pinhole throughout the day. Pointed out this is because the IFS needed to be moved under the solar beam after the wedged CaF2 window is installed. Gregor confirmed this suggestion with Axel Keens.
  • IFS seems to be holding vacuum reasonably well (<1 hPa over about a day).
  • In the evening, we handed over control of the IFS, tracker and Norhof to the herc. The tracker wasn't able to talk to the herc.
  • Getting frozen alarms with the Norhof system.

Back to top

September 29, 2016

  • Norton figured out the tracker comms problem - incorrect switch in algo
  • We had some sun in the morning, so Gregor reset the offset angles on the tracker that he was trying to use to compensate for the CaF2 wedge, and we all slowly and carefully moved the IFS under the beam. Everyone a bit happier with the tracker alignment now.
  • Orfeo working on the tracker heater and temp sensor cables.
  • Asked Stephen to pain the I-beam rails on the roof matte black to prevent saturating the webcam with reflected sunlight. May need him to put a line or two of electrical tape on the top of the enclosure as well for the same reason.
  • Norhof system still giving freeze alarms, despite changing the puck to one with a vent hold, and increasing the length of the fill tube written into EEPROM, which the manual suggests to prevent these alarms. The pump has been left outside the LN2 tank to try to dry out.
  • Applied lots of caulking inside light tube up to tracker to seal.

Outstanding Issues

Norhof system

  • Still getting freeze alarms, but after a couple of days, we need to try to put the Norhof pump back into LN2.
  • Not connected to the herc, either. We should either ask Stephen to go in every morning and push the fill button on the pump to ensure the InSb stays cold, somehow get him to connect to the Norhof system with the app and change the fill repeat time to 24 hours, or ask him to manually fill the InSb detector every day.

Hercules

  • Keeps crashing. Not sure why. We will try to disconnect items slowly until we achieve some sort of stability.
  • Can't swap removable disks without powering down the herc for the new SATA SSDs. May need to switch to a USB enclosure system.
  • Norhof system seems to communicate differently to the Caltech one. It is from 2008, and there are other indications that there are differences (e.g., you can't turn off the repeat time).

Pump

  • Pumps down well immediately after flushing and cleaning, and then shortly thereafter is unable to achieve a decent vacuum.
  • Every time it opens, there's dust in the fittings.

 

Back to top

Tag page

Files 5

FileSizeDateAttached by 
 ETL HCL ILS Test Sept 24, 2016_JM.pdf
HCl cell test Sept 24,2016 - 50 scans different pressure
2.76 MB14:48, 25 Sep 2016mendoncaActions
 GPSInfo.pdf
GPS output at ETL
81.32 kB09:46, 3 Oct 2016dwunchActions
 HCl_InSb_240916_1.dpt
HCL_2016Sept24_50scans
855.66 kB09:47, 24 Sep 2016mendoncaActions
 Security_Settings_SolarTracker_Win7_Java18.pdf
JAVA setup instructions
3.54 MB05:40, 23 Sep 2016dwunchActions
 Zeno Default Records settings in Data output menu.docx
original records from data output menu on zeno
3.47 MB15:22, 25 Sep 2016mendoncaActions